OpenStudy (anonymous):

First, rewrite 8/25 and 3/10 so that they have a common denominator. Then,< use> and = . to order 8/25 and 3/10 .

OpenStudy (anonymous):

8/25 = 32/100 3/10 = 30/100 8/25 > 3/10